[Hallc_running] [Revised Logentry] RC Update - July 20, 2025

kinney at jlab.org kinney at jlab.org
Sun Jul 20 09:30:02 EDT 2025


Logentry Text:
--
Relatively smooth production running has occurred over the past 24 hours. The z=0.36 kinematic was completed in mid-Day shift and we moved the running at z=0.9, which was completed shortly after midnight. We then set the SHMS magnet for z=0.67 point. There were some beam off periods in late swing and mid Owl, but mostly smooth running, Currents were lower than planned due to high rates in S1X; SHMS 3/4 was tolerable. During day shift will plan to have one of the 30 min runs be around 40 uA each cycle. This will bring SHMS 3/4 to about 900 kHz and S1x to about 1.7-1.8 MHz, in order to boost our statistics in a clean way, we hope.

Saturday afternoon, Dave Mack came in and we performed our first BCM calibration with max current 45 uA.

We had a coda crash around 4:45am this morning, which responded to the usual fix.

After analyzing recent data, Peter has noticed that some of the scalers are impossibly large; he will post a log entry with details soon.

The plan is to continue collecting data at z=0.56 for the next 24 hrs at least. I have requested a beam energy measurement be performed some time early in the coming week at the convenience of MCC.

Given the steady data collection, it seems we can skip having an RC meeting today (Sunday), so our next meeting will be Monday at the usual time!

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4422528
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250720/ea775896/attachment.htm>


More information about the Hallc_running mailing list